OpenStudy (anonymous):

A function g is defined for all real numbers and has the following property: g(a+b)-g(a)= 4ab+2b^2 Find g'(x)

hartnn (hartnn):

do you know the first principle definition of derivative ??

OpenStudy (anonymous):

no...

hartnn (hartnn):

\(\huge f'(x)=\lim \limits_{h \rightarrow 0}\dfrac{f(x+h)-f(x)}{h}\) haven't seen this ?

hartnn (hartnn):

thats the definition.

hartnn (hartnn):

so, for your question, first find , [g(a+b)-g(a)] /b =.... ? can you ?

OpenStudy (anonymous):

do I factor out b from the right and divided everything by b?

hartnn (hartnn):

yes! do that , and tell me what you get ?

OpenStudy (anonymous):

\[\frac{ g(a+b)-g(a) }{b }=4a+2b\]

hartnn (hartnn):

correct ! but you want, \(\huge g'(x)=\lim \limits_{h \rightarrow 0}\dfrac{g(x+h)-g(x)}{h}=....\) right ? so, just replace 'a' by 'x' and 'b' by 'h' in your last equation, what you get ?

OpenStudy (anonymous):

\[\frac{ g(x=h)-g(x) }{ h }=4x+2h??\]

hartnn (hartnn):

4x+2h is correct! \(\large g'(x)=\lim \limits_{h \rightarrow 0}\dfrac{g(x+h)-g(x)}{h}=\lim \limits_{h \rightarrow 0} (4x+2h)=.....?\) can you solve that limit ?

hartnn (hartnn):

just plug in h=0 there!

OpenStudy (anonymous):

4x?

hartnn (hartnn):

yes, \(\large g'(x)=4x\) is correct final answer :) Ask if any doubts.

OpenStudy (anonymous):

oh I get it actually thank you!

hartnn (hartnn):

welcome ^_^